# Scaling-aware rating of count forecasts

### v3

@@ 4 Application to M5-competition-models
As a first application, we use our scaling-aware rating on several models for the validation period of the M5 competition (Makridakis et al., 2022a, b; kaggle, 2022a).
@@
    Baseline. The baseline model described in Section 2.6 (details in Appendix A) is calibrated and maximally (Poisson) sharp.
@@
    Naïve-1-day-ahead. A mediocre model uses a one-day-ahead heuristic that takes yesterday’s actual sales value as prediction for today’s sales, effectively a one-day-horizon.
@@
    Global forecast. An undifferentiated forecast produces the overall average of the Simple-28-day-model for every SKU and every day.
@@
    LightGBM. The LightGBM model (kaggle, 2022b) uses an implementation of gradient boosting framework with the tree-based learning algorithm (Ke et al., 2017). It was trained with a Poisson objective and RMSE as a metric, and uses some additionally engineered features such as lagged sales. It predicts on horizon one to horizon 28, and scored an above-average performance on the public leaderboard.
To rate the models, we bucket predictions by their value, and plot the bias and the NMRPS for each bucket in Figure 7. We clip small predictions at 0.01 (which gives predictions with $r_{j}<0.01$ an advantage because they are underpredicting on average), and we clip the bias factor at 10 and 1/10 (giving the otherwise even more heavily biased prediction buckets an advantage) to fit into the visual representation. The representations allow us to judge the models and discuss their strengths and weaknesses.
@@ 4.1 Model comparison: Rate-stratified representation
In the bias scatter plot in the upper panel of Figure 7, the baseline (blue) is almost perfect by construction, and LightGBM (orange) exhibits a good level of bias for moderate predicted rates of 1 and above. For smaller rates ($r<0.1$), LightGBM-predictions are typically under-predictions, witnessing overfitting. The Naïve-1-day-ahead prediction (green) contains many 0-predictions (which were clipped to 0.01), which are often under-forecasts: When a product is not sold on a given day, the naïve-1-day-ahead-prediction for the next day is zero, but the expected value of its sales for the next day is larger than zero. This regression-to-the-mean also manifests in forecasts that follow days with abnormally many sales, these are then typically over-predictions. The green circles consequently all lie above the ideal line for predictions larger than 1 (the overall average daily selling rate is about 1.4). The Simple-28-day-model does not exhibit the same level of balance as LightGBM, but it is less biased than the naïve-1-day-ahead model. The global model produces 1.39 for every SKU, a slight underforecasting, since the overall mean sales are about 1.44.
The stratified noise plot in the lower panel of Figure 7 clearly separates the five models: The baseline fits the Poisson-ideal perfect line (reproducing Figure 2). For every prediction bucket, LightGBM is more noisy, rated in the “good” range. The Simple-28-day model is, again, slightly worse, populating the “OK” area. The scattered circles follow the parametrized corridors with $\gamma=1.5$ remarkably well – any departure of these corridors is, thus, a forecasting abnormality that should be investigated. The naïve-1-day-ahead model follows with some distance to the others on the boundary to “insufficient”. The largest two prediction buckets of that model leave the bulk – the regression to the mean is especially pronounced for the largest predictions, which correspond to the largest observations on the day before. Not surprisingly, the global model exhibits a large degree of noise, close to “unacceptable”.
@@
Model $S_{\text{overall}}$(NMRPS) Quality (NMRPS) $S_{\text{overall}}$(Bias) Quality (Bias) Bias factor
@@
With Figure 7 and the resulting ratings, one gains the confidence that the LightGBM-model is consistently better than Simple-28-day, across all prediction buckets and in a stable fashion. The rating of bias per bucket is more stringent than the overall bias alone: The Naïve-1-day-ahead method is, by construction, overall unbiased (it uses the sales of the last day before and of the first 27 days of the validation period), but only the bucket with $R\approx 1$ is unbiased, resulting in a “Fair” overall bias performance. The “insufficient” rating of the global model reflects its lack of any individualisation among SKUs and days.
@@ 4.2 Dataset comparison
The comparison of unscaled metrics and scaling-aware ratings per department allows to appreciate the added insight from the scaling-aware approach: From the WMAPE, MAE, (N)MRPS alone (displayed in Fig. 9), one cannot make any statement about which department is performing better or worse – context is required. Using the rating, and from the lower panel of Figure 8, we see that “Household 1” has less noise than “Foods 3”, and both these departments perform better than “Hobbies 2”. “Foods 3” enjoys a lack of overall bias (quotient 1.0) and an excellent bias rating (91.3%), reflecting the lack of substantial bias in every bucket. The largest predictions in “Household 1” are clearly deteriorated in both the bias and the noise plots. In “Foods 3”, only the small predictions ($r<0.3$) stand out as being under-forecasting, a general minor mis-calibration of the LightGBM model (see upper panel of Figure 7).
@@ 5 Application to Demand Edge for Retail predictions at Sainsbury’s
The M5 dataset discussion in the previous section has the great advantage to be reproducible by anyone. On the other hand, one might argue that the setup is artificial, since one deals with a curated dataset that does not contain all data that are used in production (e.g. the price of the product). To complement our discussion and exhibit the applicability of our method in practice, we apply the scaling-aware rating to the predictions of the Demand Edge for Retail forecasts produced by Blue Yonder for the UK retailer Sainsbury’s. Demand Edge for Retail is based on a causal machine-learning model that computes the contribution of features such as price, weather, day-of-the-week, holiday, promotion and many others to the expected demand of a product Wick et al. (2019, 2021). It produces a forecast that consists of a prediction for the expected mean as well as for the variance. That is, the prediction itself estimates the super-Poissonian uncertainty. To simplify the discussion and make it comparable to the M5 dataset discussion, we only use the predicted mean, and interpret it as predicted mean of the Poisson distribution.
We analyse the next-day-forecasts (horizon 1) for Monday, October 2nd to Saturday, October 28th, 2023, excluding Sundays. We exclude items that are not sold in units of pieces but volume (liters) or mass (KGs), predictions that are smaller than 0.1 / product-location-day, and unsold product-day-location combinations for which the prediction was 20 or larger. The latter filter is motivated as follows: A Poisson prediction of $\mu=20$ forecasts the probability to sell 0 items to be about $2\cdot 10^{-9}$. Thus, even if the model were grossly overestimating, a sales value of 0 is absolutely unexpected and clearly indicates an out-of-stock situation. After data cleaning, we remain with a total of more than 250 million product-location-day combinations, corresponding to more than 500 million sold items. The global overprediction of the model is less than 1%.
@@
The direct comparison of scaling-unware metrics would not be meaningful Kolassa (2008), but we can compare Figures 7 and 10 to perform a hybrid comparison: We see how the Demand Edge for Retail model outperforms the M5 competition models, since, for a given rate bucket, it produces lower values in NMRPS. This is also reflected by the excellent values $S_{\text{overall}}\text{NMRPS}=84.7\%$ and $S_{\text{overall}}\text{Bias}=78.7\%$ on the Sainsbury’s dataset.
Our approach hence confirms that, on a global level, the prediction is comparable to what could possibly be expected from first principles (the forecast is entirely product-based, since the supermarket visitors are not tracked (Tichy, 2023c)), as discussed in Section 2.1. A more in-depth investigation confirms that the reasons for non-Poissonian behavior described in Section 2.2 above are indeed relevant in this scenario (there are, for example, substantial “multi-buy” events when customers buy several pieces of one item). A thorough quantitative understanding of the mild departure from the Poisson ideal case based on those effects would be highly appreciated, but remains beyond the scope of the current work.
@@
Table 2: Metrics for Blue Yonder’s Demand Edge for Retail forecasts at Sainsbury’s. Due to non-disclosure reasons, we do not show the absolute total sales nor the sales per item-location-day, but state the sales per product-location-day normalized to the overall mean velocity. Similary to Table 1, the scaling-unaware metrics lead to ambiguous interpretations: Judging from the MAE and MRPS, one would focus to improve the Fresh assortment, based on WMAPE and NMRPS, one would be led to investigate Non foods.
Similarly to our artificial example earlier in Section 2.6, the naïve scaling trap is lurking. Large differences between the three product groups could lead an analyst to conclude a dramatically different forecasting performance. More insight can be gained via the prediction-bucketed representation in Figure 11: The upper panel reflects the low bias per prediction bucket (although predicted slow movers are slightly underestimated). The lower panel confirms that the differences in forecast quality between the product groups is small, the difference of achieved KPIs in Table 2 is mostly borne by the different selling velocities (Fresh sells at about twice the mean velocity, Non foods sells at about one third of the mean velocity). When asking for improvement potential, it is worth noting, though, that Fresh is performing slightly better than Ambient and Non foods in the slow- to mid-mover regime (mean prediction between 1/day to 10/day). The outliers in Ambient for fast movers (mean prediction $>$100/day) are worth investigating, but constitute only a tiny fraction of Ambient product-location-days and sales. Since circle size scales with the square-root of the number of measurements, small buckets are visually over-emphasized.
@@
In practice, applying scaling-aware forecast rating at Sainsbury’s has resolved many ambiguous situations in which it was not clear whether certain achieved KPIs were expected or out of range. When introducing model changes, the prediction-bucketed representation was handy to give confidence that changes have a global positive impact, across all velocities. For the forecast provider Blue Yonder, adhering to a clear methodology on how to judge models has streamlined and improved support capabilities. For the forecast consumer Sainsbury’s, the scaling-aware analyses have helped tackling data quality problems in a focused manner to improve the forecast, and given the visibility and confidence of how it is performing.
@@
6 Conclusions
Applied forecast judgement methodology should reconcile prevailing business practice (Carvallo, 2021; Bower, 2023) with the state of the art in probabilistic forecasting (Czado et al., 2009; Gneiting, 2011; Kolassa, 2016; Wei and Held, 2014; Gneiting and Katzfuss, 2014). scaling-aware forecast rating provides a framework to tackle the problem of benchmarking and precision scaling in a systematic way. The counting statistics – the scaling of the Poisson distribution and of the noise of forecast quality references – can be modeled universally, allowing a separation of these general effects and natural bounds from industry-specific aspects, captured by the parameters that define “excellent” to “unacceptable”. By adapting the parametrization introduced in Section 3.3, one can apply scaling-aware forecast rating to different forecasting problems, e.g. to industries in which bad data quality is common or certain influencing factors are known to be uncontrollable.
@@ 6.2 Future venues
Our proposed method – computing expectation values of metrics, and bucketing by similar predictions – can be applied to all probabilistic forecasts for which one can characterize the ideal case.
Within the application in retail, several generalizations are thinkable: In a business context, it is advisable to judge the forecast by its actual economic impact (Syntetos et al., 2010), which can be rated in the context of what could be possibly achieved. A certain out-of-stock or waste loss, expressed in currency, can then be given a rating (“perfect”, “excellent”, etc.). The visibility of the financial impact of forecasting error, together with the knowledge about what could be achieved if the forecast were improved to a certain quality level, will help allocate resources efficiently. The question naturally emerges to which extent human overrides of forecasts (Khosrowabadi et al., 2022) intuitively incorporate precision scaling.
The exponent $\gamma=1.5$ in the parametrization in Eq. (11) works remarkably well and quite universally, for many different applications and industries (compare the M5-analysis to the Sainsbury’s analysis – models on both datasets follow the shape of the reference lines very well, across many orders of magnitude). Nevertheless, a better theoretical explanation or empirical corroboration of that value would be appreciated. Our hope is fueled by the fact that scaling laws often reflect underlying system structures (West, 2018; Athanasopoulos and Kourentzes, 2022).
As a simple (one could even say: simplistic) first-order approximation to a benchmark, we used the Poisson distribution. For a second-order approximation, the distribution should be corrected for multibuys (a single customer buys several items of the same product), possible finite capacity and other effects. For example, the remaining gap in the bucketed representation of the Sainsbury’s dataset is, to a great extent, understandable by multibuy-effects.
@@ Acknowledgements
The authors would like to acknowledge helpful discussions with Ulla Gebbert, Sönke Niekamp, Hannes Sieling and Martin Wong, and comments on the manuscript by Martin Görner and Alexander Berkner. They are grateful to Sainsbury’s to allow us to showcase our method on their productive data.

### v4

@@ 4 Application to M5-competition-models
As a first application, we apply our scaling-aware rating to several models for the validation period of the M5 competition (Makridakis et al., 2022a, b; kaggle, 2022a).
@@
    Baseline. The baseline model described in Section 2.6 (details in A) is calibrated and maximally (Poisson) sharp.
@@
    Naïve-1-day-ahead. A mediocre model uses a one-day-ahead heuristic that takes yesterday’s observed sales value of every product-location combination as prediction for today’s sales, effectively a one-day-horizon.
@@
    Global forecast. An undifferentiated collective forecast produces the overall average of the Simple-28-day-model for every SKU and every day.
@@
    LightGBM. The LightGBM model (kaggle, 2022b) implements a tree-based learning algorithm within the gradient boosting framework (Ke et al., 2017). It was trained with a Poisson objective and RMSE as a metric, and uses some additionally engineered features such as lagged sales. It predicts on horizon one to horizon 28, and scored an above-average performance on the public leaderboard.
To rate the models, we follow the procedure described above in Section 3: We bucket predictions by their value, and plot the bias and the NMRPS for each bucket in Figure 7. We clip small predicted rates at 0.01 (which gives predictions with $r_{j}<0.01$ an advantage because they are under-predicting, on average), and we clip the bias factor at 10 and 1/10 (giving the otherwise even more heavily biased prediction buckets an advantage) to fit into the visual representation. The plots allow us to judge the models and discuss their strengths and weaknesses.
@@ 4.1 Model comparison: Rate-stratified representation
In the bias scatter plot in the upper panel of Figure 7, the baseline (blue) is almost perfect by construction, and LightGBM (orange) exhibits a good level of bias for moderate predicted rates of 1 and above. For smaller rates ($r<0.1$), LightGBM-predictions are typically under-predictions, witnessing overfitting. The Naïve-1-day-ahead prediction (green) contains many 0-predictions (which were clipped to 0.01), which are often under-forecasts: When a product is not sold on a given day, the naïve-1-day-ahead-prediction for the next day is zero, but the expected value of its sales for the next day is larger than zero. This regression-to-the-mean also manifests in forecasts that follow days with abnormally many sales, these are then typically over-predictions. The green circles consequently all lie above the ideal line for predictions larger than 1 (the overall average daily selling rate is about 1.4). The Simple-28-day-model is not as unbiased as LightGBM, but it is still less biased than the naïve-1-day-ahead model. The global model produces 1.39 for every SKU, a slight under-forecasting, since the overall mean sales are about 1.44.
The stratified noise plot in the lower panel of Figure 7 clearly separates the five models: The baseline fits the Poisson-ideal perfect line (reproducing Figure 2). For every prediction bucket, LightGBM is more noisy, rated in the “good” range. The Simple-28-day model is, again, slightly worse, populating the “OK” area. The scattered circles follow the parametrized corridors with $\gamma=1.5$ remarkably well – any departure of these corridors is, thus, a forecasting abnormality that should be investigated. The naïve-1-day-ahead model follows with some distance to the others on the boundary to “insufficient”. The largest two prediction buckets of that model leave the bulk – the regression to the mean is especially pronounced for the largest predictions, which correspond to the largest observations on the day before. Not surprisingly, the collective global model exhibits a large degree of noise, close to “unacceptable”.
@@
Model $S_{\text{overall}}$(NMRPS) Quality $S_{\text{overall}}$(Bias) Quality Bias factor
@@
With Figure 7 and the resulting ratings at hand, one gains the confidence that the LightGBM-model is consistently better than Simple-28-days, across all prediction buckets and in a stable fashion. The rating of bias per bucket is more stringent than the overall bias alone: The Naïve-1-day-ahead method is, by construction, overall unbiased (it uses the sales of the last day before and of the first 27 days of the validation period), but only the bucket with $R\approx 1$ is unbiased, resulting in a “Fair” overall bias performance. The “insufficient” rating of the global collective model reflects its lack of any individualisation among SKUs and days.
@@ 4.2 Dataset comparison
The comparison of unscaled metrics and scaling-aware ratings per department allows us to appreciate the added insight from the scaling-aware approach: From the WMAPE, MAE, (N)MRPS alone (displayed in Fig. 9), one cannot make any statement about which department is performing better or worse – context is required. Using the rating, and from the lower panel of Figure 8, we see that “Household 1” has less noise than “Foods 3”, and both these departments perform better than “Hobbies 2”. “Foods 3” enjoys a lack of overall bias (quotient 1.0) and an excellent bias rating (91.3%), reflecting the lack of substantial bias in every bucket. The largest predictions in “Household 1” are clearly deteriorated in both the bias and the noise plots. In “Foods 3”, only the small predictions ($r<0.3$) stand out as being under-forecasting, a general minor mis-calibration of the LightGBM model (see upper panel of Figure 7).
@@ 5 Application to Demand Edge for Retail predictions at Sainsbury’s
The M5 dataset discussion in the previous section has the great advantage to be reproducible by anyone. On the other hand, one might argue that the setup is artificial, since one deals with a curated dataset that does not contain all data that are used in production (e.g. the price of the product). To complement our discussion and prove the applicability of our method in practice, we apply the scaling-aware rating to the predictions of the Demand Edge for Retail forecasts produced by Blue Yonder for the UK retailer Sainsbury’s. Demand Edge for Retail is based on a causal machine-learning model that computes the contribution of features such as price, weather, day-of-the-week, holiday, promotion and many others to the expected demand of a product (Wick et al., 2019, 2021). It produces a forecast that consists of a prediction for the expected mean as well as for the variance, technical details can be found in (Wick, a, b, c). That is, the prediction itself estimates the super-Poissonian uncertainty. To simplify the discussion and make it comparable to the M5 dataset discussion, we only use the predicted mean, and interpret it as predicted mean of the Poisson distribution.
We analyse the next-day-forecasts (horizon 1) for Monday, October 2nd to Saturday, October 28th, 2023, excluding Sundays. We exclude items that are not sold in units of pieces but volume (liters) or mass (KGs), predictions that are smaller than 0.1 / product-location-day, and unsold product-day-location combinations for which the prediction was 20 or larger. The latter filter is motivated as follows: A Poisson prediction of $\mu=20$ forecasts the probability to sell 0 items to be about $2\cdot 10^{-9}$. Thus, even if the model were grossly overestimating, a sales value of 0 is absolutely unexpected and clearly indicates an out-of-stock situation. After data cleaning, we remain with a total of more than 250 million product-location-day combinations, corresponding to more than 500 million sold items. The global over-prediction of the model is less than 1%.
@@
The direct comparison of scaling-unaware metrics would not be meaningful Kolassa (2008), but we can compare Figures 7 and 10 to perform a hybrid comparison: We see how the Demand Edge for Retail model outperforms the M5 competition models in the sense that, for a given rate bucket, it produces lower values in NMRPS. This is also reflected by the excellent values $S_{\text{overall}}\text{NMRPS}=84.7\%$ and $S_{\text{overall}}\text{Bias}=78.7\%$ on the Sainsbury’s dataset.
Our approach hence confirms that, on a global level, the prediction is comparable to what could possibly be expected from first principles (Tichy, 2023c), as discussed in Section 2.1. A more in-depth investigation confirms that the reasons for non-Poissonian behavior described in Section 2.2 above are indeed relevant in this scenario (there are, for example, substantial “multi-buy” events when customers buy several pieces of one item). A thorough quantitative understanding of the mild departure from the Poisson ideal case based on those effects would be highly appreciated, but remains beyond the scope of the current work.
@@
Table 2: Metrics for Blue Yonder’s Demand Edge for Retail forecasts at Sainsbury’s. Due to non-disclosure reasons, we neither show the absolute total sales nor the sales per item-location-day, but state the sales per product-location-day normalized to the overall mean velocity. Similarly to Table 1, the scaling-unaware metrics lead to ambiguous interpretations: Judging from the MAE and MRPS, one would focus to improve the Fresh assortment, based on WMAPE and NMRPS, one would investigate Non foods.
Similarly to our artificial example earlier in Section 2.6, the naïve scaling trap is lurking. Large differences between the three product groups could lead an analyst to conclude a dramatically different forecasting performance. More insight can be gained via the prediction-bucketed representation in Figure 11: The upper panel reflects the low bias per prediction bucket (although predicted slow movers are slightly underestimated). The lower panel confirms that the differences in forecast quality between the product groups is small, the difference of achieved KPIs in Table 2 is mostly borne by the different selling velocities (Fresh sells at about twice the mean velocity, Non foods sells at about one third of the mean velocity). When asking for improvement potential, it is worth noting that Fresh is performing slightly better than Ambient and Non foods in the slow- to mid-mover regime (mean prediction between 1/day to 10/day). The outliers in Ambient for fast movers (mean prediction $>$100/day) are worth investigating, but constitute only a tiny fraction of Ambient product-location-days and sales. Since the circle size scales with the square-root of the number of measurements, small buckets are visually over-emphasized.
@@
In practice, applying scaling-aware forecast rating at Sainsbury’s has resolved many ambiguous situations in which stakeholders were unsure whether certain achieved KPIs were expected or out of range. When introducing model changes, the prediction-bucketed representation was handy to give confidence that changes have a global positive impact, across all velocities. For the forecast provider Blue Yonder, adhering to a clear methodology on how to judge models has streamlined and improved support capabilities. For the forecast consumer Sainsbury’s, the scaling-aware analyses have helped tackling data quality problems in a focused manner to improve the forecast, and given the visibility and confidence of how it is performing.
@@
6 Conclusions and outlook
Applied forecast judgement methodology should reconcile prevailing business practice (Carvallo, 2021; Bower, 2023) with the state of the art in probabilistic forecasting (Czado et al., 2009; Gneiting, 2011; Kolassa, 2016; Wei and Held, 2014; Gneiting and Katzfuss, 2014). Scaling-aware forecast rating provides a framework to tackle the problem of benchmarking and precision scaling in a systematic way. The counting statistics – the scaling of the Poisson distribution and of the noise of forecast quality references – can be modelled universally. This allows to separate these general effects and natural bounds from industry-specific aspects, which are captured by the parameters that define “excellent” to “unacceptable”. By adapting the parametrization introduced in Section 3.3, one can apply scaling-aware forecast rating to different forecasting problems, e.g. to industries in which bad data quality is common or certain influencing factors are known to be uncontrollable.
@@ 6.2 Future venues
Our proposed method – computing expectation values of metrics, and bucketing by similar predictions – can be applied to all probabilistic forecasts for which one can characterize the ideal case. Within the application in retail, several generalizations are thinkable: In a business context, it is advisable to judge the forecast by its actual economic impact (Syntetos et al., 2010), which can be rated in the context of what could be possibly achieved. A certain out-of-stock or waste loss, expressed in currency, can then be given a rating (“perfect”, “excellent”, etc.). The visibility of the financial impact of forecasting error, together with the knowledge about what could be achieved if the forecast were improved to a certain quality level, will help allocate resources efficiently. The question naturally emerges to which extent human overrides of forecasts (Khosrowabadi et al., 2022) intuitively incorporate precision scaling.
The exponent $\gamma=1.5$ in the parametrization in Eq. (11) works remarkably well and quite universally, for many different applications and industries (compare the M5-analysis to the Sainsbury’s analysis – models on both datasets follow the shape of the reference lines very well, across several orders of magnitude). Nevertheless, a better theoretical explanation or empirical corroboration of that value would be appreciated. Our hope is fuelled by the fact that scaling laws often reflect underlying system structures (West, 2018; Athanasopoulos and Kourentzes, 2022).
As a simple (one could even say: simplistic) first-order approximation to a benchmark, we used the Poisson distribution. For a second-order approximation, the distribution should be corrected for multi-buys (a single customer buys several items of the same product), possible finite capacity and other effects. The gap between empirical data and Poisson ideal in the bucketed representation of the Sainsbury’s dataset is, to a great extent, borne by multi-buy-effects.
For non-individualized retail, it is straightforward to identify the Poisson distribution as an ideal, maximally sharp distribution (Tichy, 2023c). For other processes, it might be much more involved to work out the sharpest possible forecast. Also, it is not guaranteed that the segregation by buckets of similar predicted means is sufficient: More in general, one will need to bucket in a way that “similar forecasts” (in terms of the distributions that they describe) are grouped and evaluated together. The size and number of the buckets should then ensure a good balance between precision (many small, homogenous buckets) and sufficient statistics per bucket (few large more inhomogeneous buckets). The general idea put forward in this article, segregating a large and heterogeneous group of forecasts into homogeneous buckets, is independent of the Poisson distribution.
@@ Acknowledgements
The authors would like to acknowledge helpful discussions with Ulla Gebbert, Sönke Niekamp, Hannes Sieling and Martin Wong, and comments on the manuscript by Martin Görner, Alexander Berkner and Pascal Jordan. They are grateful to Sainsbury’s to allow us to showcase our method on their productive data.